40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun May 16 12:51:08 2010, kiwamu, Update, Green Locking, reflected beam at PD 
    Reply  Sun May 16 19:25:45 2010, Koji, Update, Green Locking, reflected beam at PD 
       Reply  Thu May 20 13:29:40 2010, kiwamu, Update, Green Locking, mode profile at PPKTP crystal PPKTPmode.png
          Reply  Thu May 20 14:18:59 2010, kiwamu, Update, Green Locking, mode profile of 40m cavity MMT40mcavity.png
             Reply  Wed May 26 04:14:21 2010, kiwamu, Update, Green Locking, locked OLTF_endPDH.png
                Reply  Wed May 26 14:38:02 2010, Koji, Update, Green Locking, locked 
                   Reply  Thu May 27 02:22:24 2010, kiwamu, Update, Green Locking, more details stimes.pngspectrum.pngETMX_small.png
                      Reply  Thu May 27 09:43:50 2010, rana, Update, Green Locking, more details 
                         Reply  Mon Jul 12 13:16:53 2010, kiwamu, Update, Green Locking, PZT feedback at X end arm_day_night.png
                      Reply  Wed Jun 9 11:13:11 2010, kiwamu, Update, Green Locking, lock with PDH box 
                         Reply  Fri Jun 11 11:54:42 2010, kiwamu, Update, Green Locking, end PDH with thermal feedback LPF.pngthermal_feedback.png
                            Reply  Fri Jun 11 13:32:28 2010, Koji, Update, Green Locking, end PDH with thermal feedback 
                         Reply  Sun Jun 20 18:28:25 2010, kiwamu, Update, Green Locking, Re: lock with PDH box 
                Reply  Wed Jul 21 19:11:18 2010, Dmass, Update, Green Locking, locked 
Message ID: 2992     Entry time: Wed May 26 14:38:02 2010     In reply to: 2988     Reply to this: 2997
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: locked 

Congratulation! Probably you are right, but I could not get this is a real lock or something else.

1) How much was the fringe amplitude (DC) of the reflected beam? (Vref_max=XXX [V] and Vref_min=YYY [V])
    Does this agree with the expectation?

2) Do you have the time series? (V_ref and V_error)

Quote:

I guess I succeeded in locking of the cavity with the green beam 

 Strictly speaking, the laser frequency of the end NPRO is locked to the 40 meter arm cavity.

Pictures, some more quantitative numbers and some plots are going to be posted later.

 


After the alignment of the cavity I could see DC fringes in its reflection. Also I could see the cavity flashing on the monitor of  ETMY_CCD.

I drove the pzt of the NPRO with f=200kHz, and then the spectrum analyzer showed 200kHz beat note in the reflection signal. This means it's ready to PDH technique.

And then I made a servo loop with two SR560s, one for a filter and the other for a sum amp.

After playing with the value of the gain and the sign of the feedback signal, the laser successfully got lock. 

 

To make sure it is really locked, I measured the open loop transfer function of the PDH servo while it stayed locked. The result is shown in the attached figure.

The measured data almost agrees with the expected curve below 1kHz, so I conclude it is really locked.

However the plot looks very noisy because I could not inject a big excitation signal into the loop. If I put a big excitation, the servo was unlocked.

The current servo is obviously too naive and it only has f-1 shape, so the filter should be replaced by a dedicated PDH box as we planed.

 

ELOG V3.1.3-